# Adapting Noise to Data: Generative Flows from 1D Processes

### v3

@@ B.2 Uniform Case
Finally we want to note that the MMD functional (27) loses its convexity (along generalized geodesics) in multiple dimensions [HGBS2022], and the general existence of its Wasserstein gradient flows is unclear in the multivariate case. This yields another reason to work in 1D.
@@ Appendix C IMM with Quantile Interpolants
In this section, we want to demonstrate how the IMM framework proposed in [IMM2025] can be realized by our quantile approach. Note that in the following – for notational simplicity – we consider the one-dimensional case $X_{0},Z_{t}\in\mathbb{R}$ where we can employ quantile functions. By combining the 1D components into a multivariate model $\mathbf{X}_{0}=(X_{0}^{1},...,X_{0}^{d}),\,\mathbf{Z}_{t}=(Z_{t}^{1},...,Z_{t}^{d})$, the results of this chapter trivially extend to $\mathbb{R}^{d}$.
@@
Remark 11 (Relation to DDIM).
The interpolants used in Denoising Diffusion Implicit Models (DDIMs) [DDIM2020] are given by
@@
Proposition 12 represents the key observation which allows us to utilize our quantile process (63) in the IMM framework the same way as [IMM2025] employ the DDIM interpolants (65):
@@
The following proposition follows directly from Proposition 12 as in [IMM2025]. It is essential for deriving the appropriate loss functions.
Proposition 13.
@@
Then it is proposed in [IMM2025, Eq. (7)] to minimize the so-called naïve objective
@@
Although for marginal-preserving interpolants, a minimizer of $\mathcal{L}_{\rm naive}$ exists with minimum 0, the authors of [IMM2025] object that directly optimizing (78) faces practical difficulties when $t$ is far away from $s$. Instead, they propose to apply the following “inductive bootstrapping” technique:
@@
with a weighting function $w(s,t)$ to be chosen. The kernel $K$ of the squared MMD distance can be chosen as e.g. the (time-dependent) Laplace kernel. Importantly, the value $r$ is chosen to be a function $r=r_{s,t}\in[s,t]$ being "close to tt" and fulfilling a suitable monotonicity property.
@@
Thus, for the number of steps $n\to\infty$, it holds $0={\rm MMD}^{2}_{K}(\rho_{s},p^{\theta_{n}}(s,t_{n}))$ even for the entire interval $t_{n}\in[s,1]$. Hence, minimizing the general objective (79) with a large number of steps eventually minimizes the naïve objective (78), see [IMM2025, Theorem 1] for more details.
@@ Appendix D Radially Symmetric Processes
Definition 15 (Radial Processes).
@@
Proposition 16 (Radial Continuity).
Let $\mu_{t}$ denote the law of $\mathbf{Y}_{t}$ in Definition 15, with radially symmetric density $\rho_{t}(x)=\bar{\rho}_{t}(\|x\|)$. Let $p_{t}^{R}$ and $v_{t}^{R}$ be given as in Definition 15.
@@
By Lemma 14, it suffices to verify (86). But we have
@@
Remark 17 (Connection to Mean-Reverting Processes and Learning).
@@
The radial construction naturally extends to the quantile learning framework of Section 4. We can parametrize the radial distribution via its quantile function $Q_{R}:(0,1)\to[0,\infty)$, where $R_{t}=t\,Q_{R}(U)$ with $U\sim\mathcal{U}(0,1)$. Following Section 4.2, we learn $Q_{R}$, which can also be combined with velocity field training analogously to (41). This provides a data-adapted isotropic latent, in contrast to the fully independent coordinate-wise adaptation of our main method.
@@ Appendix E Adapting Noise to Data
E.1 Toy Target Distributions
@@
For the toy illustration in Figure 5, we work with the dataset known as Neals Funnel [neal2003slice]. The distribution of Neal’s funnel is defined as follows:
@@
$p_{\text{Checker}}(x,y)=\begin{cases}\dfrac{1}{\operatorname{area}(\mathcal{S})},&(x,y)\in\mathcal{S},\\[8.0pt] 0,&\text{otherwise}.\end{cases}$
@@
E.2 Details on the Architecture of the Learned Quantiles
We implement each one–dimensional quantile function with rational–quadratic splines (RQS) [GD1982, durkan2019neural]. We explored several ways to map $u\in(0,1)$ into the spline input; the two variants below consistently performed well and are used in our experiments. For every coordinate $i$, we write
$Q^{i}_{\phi}(u)\;=\;S^{i}_{\phi}\!\bigl(\psi(u)\bigr),\qquad u\in(0,1),$
where $S^{i}_{\phi}:\mathbb{R}\!\to\!\mathbb{R}$ is a strictly increasing RQS with an interior knot interval (B,B)(-B,B) (with $K$ bins) and linear tails outside $\pm B$ that are $C^{1}$-matched at the boundaries. The two settings differ only in the “activation” $\psi$:
Parameterization and constraints.
Per-component affine wrapper (scale/bias).
E.3 Regularization via Expected Negative Log–Jacobian
E.4 Minibatch Optimal Transport
Since the learned latent distribution is close to the data distribution, we can exploit this improved matching via an optimal transport coupling. For training, the minibatch OT is computed empirically as follows: draw a minibatch $\{\mathbf{x}_{0}^{(i)}\}_{i=1}^{B}\sim\mu_{0}$ and $\{\mathbf{u}^{(j)}\}_{j=1}^{B}\sim\mathcal{U}([0,1]^{d})$, set $\mathbf{y}^{(j)}=\mathbf{Q}_{\phi}(\mathbf{u}^{(j)})$, and define the empirical measures
@@ E.4 Minibatch Optimal Transport
The minibatch objective is
$\widehat{\mathcal{E}}_{\mathsf{Q}}(\phi)\;=\;D\big(\hat{\mu}_{0}^{B},\hat{\nu}_{\phi}^{B}\big),$
and gradients backpropagate through $\mathbf{y}^{(j)}=\mathbf{Q}_{\phi}(\mathbf{u}^{(j)})$.
Furthermore, we use the linear path $\mathbf{x}_{t}^{(j)}=(1-t_{j})\mathbf{x}_{0}^{(j)}+t_{j}\,\mathbf{y}^{(T(j))}$, $j=1,\dots,B$, with $t_{j}\sim\mathcal{U}(0,1)$, the target velocity $\mathbf{y}^{(\pi(j))}-\mathbf{x}_{0}^{(j)}$, and we optimize the empirical versions
$\widehat{\mathcal{E}}_{\mathsf{FM}}(\theta;\phi)\;=\;\frac{1}{B}\sum_{j=1}^{B}\big\|v_{\theta}\big(\mathbf{x}_{t}^{(j)},t_{j}\big)-\big(\mathbf{y}^{(T(j))}_{\phi}-\mathbf{x}_{0}^{(j)}\big)\big\|_{2}^{2},\quad\widehat{\mathcal{L}}_{\mathsf{joint}}=\widehat{\mathcal{E}}_{\mathsf{FM}}+\lambda_{Q}\,\widehat{\mathcal{E}}_{Q}.$
Appendix F Implementation Details
@@
The coupling plans are calculated using the Python Optimal Transport package [flamary2021pot]. For inference simulate the corresponding ODEs using the torchdiffeq [torchdiffeq] package. For all models we only used the batch size $128$ and learning rate 2e42e-4 for the velocities. Quantile transports are parameterised by stacked rational-quadratic splines as described in E.2, we set the minimum bin width and height to 1e31e-3 and the minimum slope to 1e51e-5.
F.1 Synthetic Examples
@@
For all models we used 3 hidden layers with width 64. We used a batch size of 128, a learning rate of 2e42e-4 and exponential moving average on the network weights of $0.999$. The baselines were trained for 200,000 iterations. We pretrain our quantiles and use the frozen quantiles during flow matching. We trained our quantile for 50,000 steps and to compensate we trained our velocity for only 100,000 steps. Note however we still train our method For the RQS we choose logit activation, 32 bins, a bound of 500 and one layer.
@@
The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 25,000. For both datasets, we trained the velocity model with 3 layers and a hidden width of 64 for 100,000 steps. For the RQS we choose the parameters number of bins 32, bound 50, layers 3.
Figure 16: Flow Matching with optimal coupling using Gaussian noise (left) and our learned noise (right) after 20k training steps with identical parameters. Generated samples are shown in blue, and ground-truth samples in red
F.2 Image Experiments
For both image datasets, we adapt the U-Net from [dhariwal2021diffusion] to parametrize our velocity field.
@@ F.2 Image Experiments
For the MNIST dataset we use the U-Net with base width 64, channel multipliers $(1,2,4)$, two residual blocks per resolution, attention at $7\times 7$, 1 attention head, and dropout 0.1. We clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.99$. The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 30,000.
@@
Here we use the U-Net with base width 128, channel multipliers $(1,2,2,2)$, two residual blocks per resolution, attention at $16\times 16$, four attention heads , and dropout 0.1. We clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.9999$. To evaluate our results, we use the Fréchet inception distance (FID) [HRUNH2017]. The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 25,000.

### v4

@@ Appendix C IMM with Quantile Interpolants
In this section, we want to demonstrate how the IMM framework proposed in (Zhou et al., 2025) can be realized by our quantile approach.
The general idea of consistency models is to predict the jumps from a process $Z_{t}$ to the target $X_{0}$, while factoring in the consistency of the trajectory of $Z_{t}$ via $Z_{s}$, $0<s<t$. In FM, this consistency of the flow is usually neglected as only single points on the FM paths are sampled. Also, consistency models as one-step or multistep samplers usually are in no need of velocity fields.
Note that in the following – for notational simplicity – we consider the one-dimensional case $X_{0},Z_{t}\in\mathbb{R}$ where we can employ quantile functions. By combining the 1D components into a multivariate model $\mathbf{X}_{0}=(X_{0}^{1},...,X_{0}^{d}),\,\mathbf{Z}_{t}=(Z_{t}^{1},...,Z_{t}^{d})$, the results of this chapter trivially extend to $\mathbb{R}^{d}$.
@@
Remark C.1 (Relation to DDIM).
The interpolants used in Denoising Diffusion Implicit Models (DDIMs) (Song et al., 2020) are given by
@@
Proposition C.2 represents the key observation which allows us to utilize our quantile process (74) in the IMM framework the same way as (Zhou et al., 2025) employ the DDIM interpolants (76):
@@
The following proposition follows directly from Proposition C.2 as in (Zhou et al., 2025). It is essential for deriving the appropriate loss functions.
Proposition C.3.
@@
Then it is proposed in (Zhou et al., 2025) to minimize the so-called naïve objective
@@
Although for marginal-preserving interpolants, a minimizer of $\mathcal{L}_{\rm naive}$ exists with minimum 0, the authors of (Zhou et al., 2025) object that directly optimizing (89) faces practical difficulties when $t$ is far away from $s$. Instead, they propose to apply the following “inductive bootstrapping” technique:
@@
with a weighting function $w(s,t)$ to be chosen. The kernel K{K} of the squared MMD distance can be chosen as e.g. the (time-dependent) Laplace kernel. Importantly, the value $r$ is chosen to be a function $r=r_{s,t}\in[s,t]$ being ”close to tt” and fulfilling a suitable monotonicity property.
@@
Thus, for the number of steps $n\to\infty$, it holds $0={\rm MMD}^{2}_{K}(\rho_{s},p^{\theta_{n}}(s,t_{n}))$ even for the entire interval $t_{n}\in[s,1]$. Hence, minimizing the general objective (90) with a large number of steps eventually minimizes the naïve objective (89), see Theorem 1 (Zhou et al., 2025) for more details.
@@ Appendix D Radially Symmetric Processes
Definition D.2 (Radial Processes).
@@
Proposition D.3 (Radial Continuity).
Let $\mu_{t}$ denote the law of $\mathbf{Y}_{t}$ in Definition D.2, with radially symmetric density $\rho_{t}(x)=\bar{\rho}_{t}(\|x\|)$. Let $p_{t}^{R}$ and $v_{t}^{R}$ be given as in Definition D.2.
@@
By Lemma D.1, it suffices to verify (97). But we have
@@
Remark D.4 (Connection to Mean-Reverting Processes and Learning).
@@
The radial construction naturally extends to the quantile learning framework of Section 4. We can parametrize the radial distribution via its quantile function $Q_{R}:(0,1)\to[0,\infty)$, where $R_{t}=t\,Q_{R}(U)$ with $U\sim\mathcal{U}(0,1)$. Following Section 5, we learn $Q_{R}$, which can also be combined with velocity field training analogously to (24). This provides a data-adapted isotropic latent, in contrast to the fully independent coordinate-wise adaptation of our main method.
@@ Appendix E Adapting Noise to Data
E.1 Counterexample: Marginal Product
For the measure
one has $W_{2}^{2}(\mu,\mu_{\mathrm{marg}})=2$, whereas for
E.2 Details on the Architecture of the Learned Quantiles
We implement each one–dimensional quantile function with rational–quadratic splines (RQS) (Gregory and Delbourgo, 1982; Durkan et al., 2019). We explored several ways to map $u\in(0,1)$ into the spline input; the two variants below consistently performed well and are used in our experiments. For every coordinate $i$, we write
$Q^{i}_{\phi}(u)\;=\;S^{i}_{\phi}\!\bigl(\psi(u)\bigr),\qquad u\in(0,1),$
where $S^{i}_{\phi}:\mathbb{R}\!\to\!\mathbb{R}$ is a strictly increasing RQS with an interior knot interval (B,B)(-B,B) (with $K$ bins) and linear tails outside $\pm B$ that are $C^{1}$-matched at the boundaries. The two settings differ only in the ”activation” $\psi$:
Parameterization and Constraints.
Per-Component Affine Wrapper (Scale/Bias).
Regularization via Expected Negative Log–Jacobian
Computational Efficiency and Scalability.
The quantile architecture is highly efficient in both computation and memory. Each component $i$ requires only $\mathcal{O}(K)$ parameters for the RQS (where $K$ is the number of bins) plus two affine parameters, totaling roughly $4K+2$ parameters per dimension for typical implementations. For a $d$-dimensional problem, this yields $\mathcal{O}(d\cdot K)$ total parameters—negligible compared to modern UNet architectures which often contain millions of parameters. Forward evaluation of $Q_{\phi}(u)$ involves $d$ independent spline evaluations operating in parallel. The diagonal Jacobian structure means that both the determinant and its gradient reduce to $d$ independent scalar derivatives with analytical closed-form expressions which are fully parallelizable, avoiding expensive automatic differentiation of matrix operations.
In practice, as noted in Section 6, the computational overhead (on CIFAR10) during joint training is approximately $2.7\%$ and drops to $0.5\%$ after freezing the quantile. Furthermore we only used 300k parameters for the quantile in contrast to 35M for the U-Net, making the approach highly scalable to high-dimensional problems. The strict monotonicity constraints and bounded parameterization (via softplus and normalization) ensure numerical stability throughout training, and we observed no instabilities across our experiments spanning dimensions from $d=2$ to $d=3072$ (CIFAR-10).
Appendix F Experiment Details
F.1 Toy Target Distributions
@@
For the toy illustration in Figure 2, we work with the dataset known as Neals Funnel (Neal, 2003). The distribution of Neal’s funnel is defined as follows:
@@
$p_{\text{Checker}}(x,y)=\begin{cases}\dfrac{1}{\operatorname{area}(\mathcal{S})},&(x,y)\in\mathcal{S},\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
@@
F.2 Loss Implementation
For training, the minibatch OT is computed empirically as follows: draw a minibatch $\{\mathbf{x}_{0}^{(i)}\}_{i=1}^{B}\sim\mu_{0}$ and $\{\mathbf{u}^{(j)}\}_{j=1}^{B}\sim\mathcal{U}([0,1]^{d})$, set $\mathbf{y}^{(j)}=\mathbf{Q}_{\phi}(\mathbf{u}^{(j)})$, and define the empirical measures
@@ F.2 Loss Implementation
The minibatch quantile alignment objective is
$\widehat{\mathcal{L}}_{\operatorname{AN}}(\phi)\;=\;W_{2}^{2}\big(\hat{\mu}_{0}^{B},\hat{\nu}_{\phi}^{B}\big),$
and gradients backpropagate through $\mathbf{y}^{(j)}=\mathbf{Q}_{\phi}(\mathbf{u}^{(j)})$. Let $T:\{1,\ldots,B\}\to\{1,\ldots,B\}$ denote the optimal assignment that minimizes $\sum_{i=1}^{B}\|\mathbf{x}_{0}^{(i)}-\mathbf{y}^{(T(i))}\|_{2}^{2}$, and define its inverse $P(j)=i$ such that $T(i)=j$. We use the conditional flow path $\mathbf{x}_{t}^{(j)}=(1-t_{j})\mathbf{x}_{0}^{(P(j))}+t_{j}\,\mathbf{y}^{(j)}$, $j=1,\dots,B$, with $t_{j}\sim\mathcal{U}(0,1)$. The target velocity is $\mathbf{y}^{(j)}-\mathbf{x}_{0}^{(P(j))}$, we apply a stop-gradient operator $\operatorname{sg}(\cdot)$ to this target in the flow matching loss. This prevents gradients from the velocity model from flowing back through the quantile function in this term, ensuring that $\mathbf{Q}_{\phi}$ is updated primarily through $\widehat{\mathcal{L}}_{\operatorname{AN}}$, while $v^{\theta}$ learns to match the transport directions defined by the current quantile map. Note however the stop gradient operation only slightly stabilizes training, we can train with full gradients as well. We optimize the empirical version
${\widehat{\mathcal{L}}_{\mathsf{CFM}}(\theta,\phi)}$ ${\;=\;\frac{1}{B}\sum_{j=1}^{B}\big\|v^{\theta}\big(\mathbf{x}_{t}^{(j)},t_{j}\big)-\operatorname{sg}\big(\mathbf{y}^{(j)}-\mathbf{x}_{0}^{(P(j))}\big)\big\|_{2}^{2}},\quad\widehat{\mathcal{L}}(\theta,\phi)=\widehat{\mathcal{L}}_{\mathsf{CFM}}(\theta,\phi)+\lambda\,\widehat{\mathcal{L}}_{\operatorname{AN}}(\phi).$
F.3 Implementation Details
@@
The coupling plans are calculated using the Python Optimal Transport package (Flamary et al., 2021). For inference simulate the corresponding ODEs using the torchdiffeq (Chen, 2018) package. For all models we only used the batch size $128$ and learning rate 2e42e-4 for the velocities. We use Adam (Kingma and Ba, 2015) as the optimizer. The quantiles are parameterised by rational-quadratic splines as described in E.2, we set the minimum bin width and height to 1e31e-3 and the minimum slope to 1e51e-5. We could in principle stack multiple RQS layers, however for all of our experiments we use one layer.
F.4 Experiment Setup: Synthetic Examples
@@
For all models we used 3 hidden layers with width 64. We used a batch size of 128, a learning rate of 2e42e-4 and exponential moving average on the network weights of $0.999$. The baselines were trained for 200,000 iterations. Since there is a very high variance when sampling from the funnel, we pretrain our quantiles and use the frozen quantiles during flow matching. We trained our quantile for 50,000 steps and to compensate we trained our velocity for only 100,000 steps. For the RQS we chose logit activation, 32 bins and a bound of 500.
@@
The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 25,000. For both datasets, we trained the velocity model with 3 layers and a hidden width of 256 for 100,000 steps. Furthermore we used sinusoidal positional embeddings for the checkerboard. For the RQS we chose logit activation, 32 bins with a bound of 50.
F.5 Experiment Setup: Image Experiments
For both image datasets, we adapt the U-Net from (Nichol and Dhariwal, 2021) to parametrize our velocity field.
@@ F.5 Experiment Setup: Image Experiments
For the MNIST dataset we use the U-Net with channel multipliers $(1,2,4)$, two residual blocks per resolution, attention at $7\times 7$, and 1 attention head. We clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.99$. We test three configurations with base widths of 8, 16, and 32 channels. For these ablation runs, we use quantile loss weight $\lambda=1.0$, regularization parameter $\beta=0.1$, and rational quadratic spline with bounded activation, 16 bins and bound 3.0. The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 over the next 10,000 steps. The images in Figure 5 were generated using our 32 channel configuration.
@@
We use exactly the same U-Net setup from (Tong et al., 2024). We clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.9999$. To evaluate our results, we use the Fréchet inception distance (FID) (Heusel et al., 2017). The quantiles were trained for the first 50,000 steps, after which the learning rate was linearly decayed to 0 by step 55,000. We used $\lambda=1$ and varied $\beta$. For the RQS we used logit activation, 32 bins and a bound of 25.
@@
Appendix G Further Experimental Results
In the following, we highlight additional experimental results that are not included in the main paper.
G.1 Synthetic Datasets
Figure 12: Samples (1M) from our learned latent of the funnel distribution. Color shows norm of the endpoint sample after solving the reverse ODE.
For the funnel experiment in Section 6.2, we visualize the learned latent distribution $\mathbf{Q}_{\phi}$ in Figure 12. We draw one million samples to accurately depict the distribution, and color each sample by the norm of the corresponding target sample after solving the ODE. The learned latent exhibits substantially heavier tails than a Gaussian, yielding improved tail behavior compared to a standard Gaussian, as also illustrated in Figure 2.
For the checkerboard distribution (see Section F.1), the velocity-field network converges in substantially fewer training iterations. After 20k training steps, the generated samples match the target distribution much more closely, see Figure  13.
In figure  14 we consider what happens when training the latent purely using the contribution from the FM loss, i.e setting $\lambda=0$.
Figure 13: Flow Matching with optimal coupling using Gaussian noise (left) and our learned noise (right) after 20k training steps with identical parameters. Starting in the learned noise the model produces much better samples. Generated samples are shown in blue, and ground-truth samples in red.
(a) Checkerboard (cf. Figure 1). Left: $\lambda=50$, Right: $\lambda=0$.
(b) Gaussian mixture (cf. Figure 3). Left: $\lambda=50$, Right: $\lambda=0$.
Figure 14: Effect of the regularization weight $\lambda$ on learned latent distributions for the checkerboard and GMM target distributions.
G.2 Image Data Set: MNIST
For MNIST, Figure 15 visualizes the learned latent distribution $\mathbf{Q}_{\phi}$ at different pixel locations (e.g., the top-left corner, the center, and the mid-right region). The learned latent matches the correct mean, while the increased variance in the top-right region is induced by the differential-entropy regularization (see Section 6.3). We again plot FID over training steps for the capacity-constrained networks, and additionally report the corresponding parameter counts in Figure 16.
Figure 16: Ablation study over capacity of the U-Net for sampling from the MNIST dataset. The FID curves show that our method achieves significantly lower FIDs for lower capacities. Note the difference in parameters is approximately 4040k.
G.3 Image Dataset: CIFAR10
In Figure 17, we present the full ablation over the regularization parameter (see Section 6.3). Most regularization settings outperform the standard Gaussian noise distribution.
Figure 18 summarizes the training dynamics over iterations, including the Wasserstein distance, the differential-entropy regularization term, and the gradient norms of both the quantile and velocity-field networks.
$\beta$ FID (20 steps) FID (100 steps)
0.1 8.93±0.04 5.22±0.02
0.2 7.81±0.04 4.75±0.02
0.3 7.48±0.05 4.53±0.05
0.4 7.60±0.05 4.54±0.01
0.5 7.66±0.03 4.49±0.02
0.6 7.70±0.05 4.47±0.03
0.7 7.93±0.05 4.59±0.02
0.8 7.77±0.05 4.42±0.02
0.9 8.09±0.04 4.56±0.02
1.0 8.35±0.03 4.66±0.04
1.1 8.60±0.04 4.80±0.03
Baseline 8.42±0.07 4.63±0.05
Figure 17: Complete FID scores on CIFAR-10 for all $\beta$ values. Our method reached the best validation FID after 320k steps, while the baseline took 340k. We used those checkpoints for the evaluation. We evaluated the FID using 5 seeds and report the mean as well as the standard deviation. Red denotes 20 step FID, blue 100 step FID, dotted line refers to baseline.
(f) Final performance metrics at 55k training steps vs. regularization parameter $\beta$.
(b) Wasserstein distance evolution during training for different $\beta$.
(c) Regularization loss $\mathbb{E}[\log|\det J_{Q}|]$ across $\beta$ values.
(d) Gradient norm of the quantile function during training.
(e) Gradient norm of the velocity field for fixed $\beta=0.5$ over training.
Figure 18: Ablation studies showing the effect of regularization and model capacity on training dynamics and final performance.